# LLMs as Narcissistic Evaluators: When Ego Inflates Evaluation Scores

### v3

@@ Abstract
Automatic evaluation of generated textual content presents an ongoing challenge within the field of NLP. Given the impressive capabilities of modern language models (LMs) across diverse NLP tasks, there is a growing trend to employ these models in creating innovative evaluation metrics for automated assessment of generation tasks. This paper investigates a pivotal question: Do language model-driven evaluation metrics inherently exhibit bias favoring texts generated by the same underlying language model? Specifically, we assess whether prominent LM-based evaluation metrics (e.g. BARTScore, T5Score, and GPTScore) demonstrate a favorable bias toward their respective underlying LMs in the context of summarization tasks. Our findings unveil a latent bias, particularly pronounced when such evaluation metrics are used in an reference-free manner without leveraging gold summaries. These results underscore that assessments provided by generative evaluation models can be influenced by factors beyond the inherent text quality, highlighting the necessity of developing more dependable evaluation protocols in the future.
@@
[…]
@@ 1 Introduction
Evaluation is a fundamental element in both tracking progress and ensuring meaningful advancements across various dimensions within the field of Natural Language Processing (NLP). Therefore, the reliability of evaluation metrics plays a critical role in this process. Evaluating generated texts is one of the challenging and open problems in NLP given that different forms can convey the same meaning. This challenge has led to the development of various evaluation metrics for tasks involving Natural Language Generation (NLG). While human evaluation by experts stands as the most reliable approach for assessing generated outputs, it is costly and time-consuming, limiting its broader use. As a result, automatic evaluation metrics have emerged as practical alternatives to keep pace with the rapid progress in NLP van der Lee et al. (2019). Recent evaluation metrics for generation tasks, such as BERTScore Zhang et al. (2020), BARTScore Yuan et al. (2021), T5Score Qin et al. (2022), GPTScore Fu et al. (2023), and G-Eval Liu et al. (2023), increasingly rely on pretrained language models.
@@
We conducted our analysis using the CNN/Daily Mail Hermann et al. (2015); Nallapati et al. (2016) and XSUM Narayan et al. (2018) summarization datasets. The assessment covers two settings: reference-based, using gold summaries for evaluation (a common approach in supervised summarization), and reference-free, comparing generated summaries against source documents (a common approach in both unsupervised summarization and factuality assessment).
Based on our analysis, we have derived the following findings: (1) Generative evaluators tend to assign higher scores to the content generated by the same underlying model. This bias becomes more pronounced when the fine-tuning configuration and model size match for both the generator and evaluator. (2) Inflated scores are particularly noticeable in the reference-free setting, which is concerning due to the popularity of this evaluation approach for assessing the factual correctness of generated texts (Koh et al., 2022). (3) Apart from the same model bias, inflated scores are also influenced by the preference for longer summaries by certain evaluators.
@@ 2 Related Work
Reference-based metrics are commonly used to evaluate text generation tasks, including summarization, by measuring the similarity between generated and reference texts. Traditionally, metrics like BLEU Papineni et al. (2002) and ROUGE Lin (2004) were employed to assess a generated text based on surface-level similarities, measured through the n-gram overlap between the generated and reference texts. However, these traditional metrics often exhibit weak correlations with human judgments, particularly when evaluating advanced generation systems Peyrard (2019); Mathur et al. (2020).
Recognizing these limitations, recent trends in summarization evaluation lean towards semantic-level assessments, moving beyond direct word overlap comparisons. Notable metrics embracing this approach include BERTScore Zhang et al. (2020), MoverScore Zhao et al. (2019), BARTScore Yuan et al. (2021), BLEURT Sellam et al. (2020), and variations thereof. By leveraging pretrained language models, these metrics focus on capturing semantic content, providing a more nuanced and accurate evaluation of summarization system outputs.
@@
With the widespread use of generation models across diverse domains, the need for reference-free evaluation metrics has surged. In response to this challenge, recent attention has been directed towards metrics that enable the evaluation of generated texts solely based on source documents, especially when annotated reference texts may not be available for new domains Böhm et al. (2019); Gao et al. (2020); Chen et al. (2021); Wu et al. (2020); Scialom et al. (2021); Honovich et al. (2021); Zhong et al. (2022); Liu et al. (2023).
@@
Despite their widespread use, automatic evaluation metrics have notable shortcomings. These metrics may not be robust when faced with challenges such as spurious correlations, noise, or out-of-domain texts (Vu et al., 2022; Sai et al., 2021; He et al., 2023; Durmus et al., 2022). Furthermore, their effectiveness diminishes when evaluating very long documents (Amplayo et al., 2022). There is also evidence suggesting a potential bias towards ranking extractive summaries higher than abstractive ones Amplayo et al. (2022).
Traditional reference-based evaluation systems, such as ROUGE or BLEU, have been criticized for their inability to measure content quality or capture syntactic errors Reiter and Belz (2009). Consequently, these metrics may inaccurately reflect the reliability of generation systems. For example, they might assign high scores to outputs that are fluent but meaningless and unfaithful, as long as many of the same words are used Gehrmann et al. (2021). Although embedding-based metrics (e.g., BERTScore) show improved performance in similarity measurement, they are still inadequate for assessing the extent of shared information between two summaries, a crucial indicator of summary information quality Deutsch and Roth (2021).
Reference-free metrics, on the other hand, exhibit a bias towards outputs generated by models that are more similar to their own Deutsch et al. (2022). To the best of our knowledge, this study represents the initial attempt to perform a systematic exploration is yet to be undertaken. Additionally, question-answering-based reference-free metrics for summarization evaluation are prone to localizing errors within summaries Kamoi et al. (2023).
Metrics based on Large Language Models, which are capable of conducting both reference-based and reference-free evaluations, typically demonstrate superior correlations with human quality judgments across diverse NLG tasks and evaluation dimensions Deutsch et al. (2022). While prior work has reported that LLM-based metrics prefer LLM-generated text, raising a concern about the shortcomings of LLMs as evaluatiorsLiu et al. (2023), our work conducts a systematic evaluation to address a fundamental question: Do language model-driven evaluation metrics inherently display bias favoring texts generated by the same underlying language model? We explore this question across both reference-based and reference-free evaluations and for a range of different Large Language Models.
@@ 3 Methodology
To investigate the impact of the same model bias—determining whether a language model-based evaluator favors outputs generated by a similar language model—we conduct a comprehensive series of experiments involving both quantitative comparisons and qualitative analysis. Our quantitative comparisons involve using language models of varying sizes and finetuning configurations as both the evaluator and generator models. This structured approach enables us to systematically examine the potential bias across different LM configurations. Subsequently, we verify the results through qualitative analysis using a subset of models’ summaries that are accompanied by human evaluation to further demonstrate that higher scores produced by evaluators do not necessarily correlate with higher quality generated outputs.
@@ 3.1 Evaluators
We describe the generation progress as follows: given a source text s, and optional human written reference r , generate a hypothesis h, which can be represented as:
@@
where h denotes hypothesis, a refers to the aspect to evaluate, and $\mathcal{S}$ denotes supplementary text (i.e., s or r) that is employed alongside evaluations in various settings Fu et al. (2023). For instance, it could be source text s in a reference-free scenario which assesses the summary based on the source article directly Fabbri et al. (2021). Whereas in the reference-based paradigm, the evaluation considers semantic overlap between the generated and reference summaries r Bhandari et al. (2020).
The evaluators utilised in our study all share a conditional probability paradigm, which can be generally concluded as the definition in:
$\!\!\!\!\!\!\!\!\mathrm{Score}(\textbf{{h}}|d,a,\mathcal{S})=\sum_{t=1}^{m}w_{t}\log{p}(h_{t}|\textbf{{h}}_{<t},\mathcal{S},\theta),$ (2)
Under a sequence-to-sequence model parameterised by $\theta$, where $d$ refers to the task description and $w_{t}$ denotes the weight of the token at position $t$. Previous works normally treat each token equally Yuan et al. (2021); Fu et al. (2023); Finally, $\mathcal{S}$ is determined according to the aspect to evaluate.
@@
Leveraging generative models to conduct evaluation has been FURTHER ADVANCED with various of more recent large language models Fu et al. (2023), showing a great performance and covering a rich variety of aspects for comprehensive evaluations. With an understanding of natural language instructions, GPTScore (including GPT-X and FLAN-T5 models) can perform intricate and personalised assessments without requiring additional training.
@@
We additionally include Cohere, the more recent language model to enrich our experiment. The evaluation scores assigned by the model is calculated according to Eq. 2, aligned with BARTScore, T5Score, and GPTScore.
@@ 3.2 Generation Models
To ensure the reproducibility of our analysis, we exclusively employ publicly available checkpoints for the utilized models. Apart from the GPT3-Curie model that is taken from the OpenAI API [^3] Note that the newer models offered by the OpenAI API, such as GPT-4 , no longer support providing token probabilities, and hence they cannot be tested in our experiment. and generation model obtained from Cohere, the rest of the models are taken from the Hugging Face model hub [^4] https://huggingface.co/models..
@@ 3.3 Datasets
The SummEval benchmark contains summaries generated by various summarization models for 100 articles from the CNN/DM test set, with each summary supplemented by human annotations. More specifically, SummEval incorporates human annotations by both expert and crowd-sourced human annotators, targeting dimensions of coherence, consistency, fluency, and relevance. Ratings are on a scale of 0 to 5, with higher values indicating better performance.
Similarly, RoSE contains summaries generated by recent generative models based on CNN/DM documents, accompanied by their corresponding human evaluations. We use 100 summaries from each of the BART and GPT-3 models from the ROSE benchmark. The RoSE benchmark proposed an assessment protocol termed ”Atomic Content Units” (ACUs). ACU score gauges quality of evaluated summaries based on whether the presence of single facts (i.e., atomic facts) from reference are included in the evaluated summaries. ACU score is calculated by ACU matching:
@@
where $\mathcal{A}$ is a set of ACUs from gold summaries and s is the candidate summary.
@@
SummEval-BART 5.00 2.67 4.56 4.67
SummEval-T5 5.00 2.33 4.46 4.67
SummEval-GPT2 5.00 1.33 3.52 3.33
Table 1: Distribution of human annotation scores on RoSE and SummEval datasets, where RoSE mainly considers the index ’ACU’ and SummEval focuses on ’Coherence’ and ’Consistency’ aspects according to expert annotators.
@@
Figure 2: Assessing Bias in the CNN/DM Dataset using heatmaps in the reference-free setting. Observing the darkest cells along the diagonal line, from the top left to the bottom right, indicates a distinct bias among evaluators towards their respective models. All evaluator scores are normalized to a range between 0 and 1. NB: the number in the bracket represents the average length of summaries (measured in words) produced by the respective model.
@@ 3.4 Quantitative Comparisons
Due to the nature of log probabilities, original scores from each evaluator should be negative, and a higher score indicates better quality according to the evaluator. When weights $w_{t}$ in Eq. 2 are treated equally, the evaluation protocols of BARTScore, T5Score, and GPTScore are all conditional probability paradigms. To ensure comparability among the scores provided by 20 distinguished evaluators, a uniform normalization process is applied to the scores generated by each evaluator. The normalization procedure standardizes the scores across a scale ranging from 0 to $\alpha$  [^7] In our work, we set parameter $\alpha$ to 1. as formulated in Eq. 4, where $S_{i,j}$ indicates scores evaluated by the $j$-th evaluator on summaries generated by the $i$-th generator.
$S_{i,j}^{norm}=\frac{\alpha(S_{i,j}-\min_{i}S_{i,j})}{\max_{i}S_{i,j}-\min_{i}S_{i,j}}$ (4)
In this context, a normalized score of $\alpha$ signifies the highest quality attributed by the generator, while a score of 0 indicates the lowest quality.
@@ 3.5 Qualitative Analysis
For qualitative analysis, we employ Spearman Correlations Zar (2014) and Kendall Correlations Freedman et al. (2007), which respectively assess monotonic relationships and order associations between human evaluations and LM evaluator scores.
For the SummEval dataset, we calculate the average value of expert annotations for coherence and consistency aspects, aligned with the reference-free input setting in the evaluation protocol specified in Yuan et al. (2021). For the evaluations based on the RoSE benchmark, ACU annotations are suited for reference-based summary salience evaluation. Therefore, we employ the correlation values obtained from the SummEval dataset for the reference-free setting and those from the RoSE benchmark for the reference-based setting.
@@ 4 Experimental Results
4.1 Quantitative Comparisons: Assessing Bias in LM-Evaluators Towards Their Own Output
Figure 3: Assessing Bias on CNN/DM Dataset using heatmaps in the reference-based setting. Observing darker cells along the diagonal line indicates potential bias. All evaluator scores are normalized to a range between 0 and 1. NB: the number in the bracket represents the average length of summaries (measured in words) produced by the respective model.
Figures 2 and 3 display heatmaps presenting evaluator scores for various summaries generated by different generators from CNN/DM documents. These scores are computed by averaging the individual scores of the selected 500 documents. In both heatmaps, we observe darker cells along the diagonal line, running from the top left to the bottom right. This indicates potential evaluator bias towards their corresponding generator models. However, this bias is notably more pronounced in the reference-free setting, commonly used for factuality evaluation Koh et al. (2022). In this context, we highlight the presence of the darkest cells along the diagonal line.
Furthermore, as shown in Figure 2, we note a distinct trend: T5-based generators, whether fine-tuned or not, tend to receive higher scores when assessed using different T5Score variations compared to evaluations using BARTScore, GPTScore, or Cohere. This results in a concentrated dark rectangle at the heatmap’s centre. Similarly, we observe a parallel trend for BART-based generators, whether fine-tuned or not. It is also quite evident that when using FLAN-T5 variants and Cohere as evaluators, they tend to favor the summaries generated by their respective models. Meanwhile, evaluators are prone to assigning higher ranks to generators under the same dataset compared to those fine-tuned on others. For instance, when utilizing T5 fine-tuned on the XSUM dataset as evaluators, a preference is observed for BART-XSUM generators over T5-vanilla models. We observe the same pattern on summaries generated based XSUM documents. The corresponding heatmap is included in Figure 6 in Appendix B.1.
It can also be observed that BART-based evaluators tend to rank FLAN-T5 generators the lowest, resulting in a normalized score of 0. This is likely to be attributed to the difference between BART and FLAN-T5’s architectures and BART-based models’ preference towards longer summaries. Similar observations can be observed that T5-based models score the summaries generated by Cohere low, as they tend to favor shorter summaries. Detailed analysis of the correlation between the length of generated summaries and the scores assigned by each evaluator is given in Figure 5.
Figure 4: Heatmaps of evaluation scores on the RoSE & SummEval benchmarks for the reference-free and reference-based setting. It is recommended to use the reference-free setting for SummEval and the reference-based setting for RoSE, aligning with the specific aspects each benchmark emphasizes.

### v4

@@ Abstract
Automatic evaluation of generated textual content presents an ongoing challenge within the field of NLP. Given the impressive capabilities of modern language models (LMs) across diverse NLP tasks, there is a growing trend to employ these models in creating innovative evaluation metrics for automated assessment of generation tasks. This paper investigates a pivotal question: Do language model-driven evaluation metrics inherently exhibit bias favoring texts generated by the same underlying language model? Specifically, we assess whether prominent LM-based evaluation metrics (e.g. BARTScore, T5Score, and GPTScore) demonstrate a favorable bias toward their respective underlying LMs in the context of summarization tasks. Our findings unveil a latent bias, particularly pronounced when such evaluation metrics are used in a reference-free manner without leveraging gold summaries. These results underscore that assessments provided by generative evaluation models can be influenced by factors beyond the inherent text quality, highlighting the necessity of developing more reliable evaluation protocols in the future.
@@
Yiqi Liu1 Nafise Sadat Moosavi2 Chenghua Lin1 1University of Manchester 2University of Sheffield yiqi.liu-6@postgrad.manchester.ac.uk n.s.moosavi@sheffield.ac.uk chenghua.lin@manchester.ac.uk
@@ 1 Introduction
Evaluation is a fundamental element in both tracking progress and ensuring meaningful advancements across various dimensions within the field of Natural Language Processing. Therefore, the reliability of evaluation metrics plays a critical role in this process. Evaluating generated texts is one of the challenging and open problems in NLP given that different forms can convey the same meaning. This challenge has led to the development of various evaluation metrics for tasks involving Natural Language Generation (NLG). While human evaluation by experts stands as the most reliable approach for assessing generated outputs, it is costly and time-consuming, limiting its broader use. As a result, automatic evaluation metrics have emerged as practical alternatives to keep pace with the rapid progress in NLP van der Lee et al. (2019). Recent evaluation metrics for generation tasks, such as BERTScore Zhang et al. (2020), BARTScore Yuan et al. (2021), T5Score Qin et al. (2022), GPTScore Fu et al. (2023), and G-Eval Liu et al. (2023), increasingly rely on pretrained language models.
@@
We conducted our analysis using the CNN/Daily Mail Hermann et al. (2015) and XSUM Narayan et al. (2018) summarization datasets. The assessment covers two settings: reference-based, using gold summaries for evaluation (a common approach in supervised summarization), and reference-free, comparing generated summaries against source documents (a common approach in both unsupervised summarization and factuality assessment).
Based on our analysis, we have derived the following findings: (1) Generative evaluators tend to assign higher scores to the content generated by the same underlying model. This bias becomes more pronounced when the fine-tuning configuration and model size match for both the generator and evaluator. (2) Inflated scores are particularly noticeable in the reference-free setting, which is concerning due to the popularity of this evaluation approach for assessing the factual correctness of generated texts (Koh et al., 2022). (3) Apart from self-bias, inflated scores are also influenced by the preference for longer summaries by certain evaluators.
@@ 2 Related Work
Reference-based metrics are commonly used to evaluate text generation tasks, including summarization, by measuring the similarity between generated and reference texts. Traditionally, metrics like BLEU Papineni et al. (2002) and ROUGE Lin (2004) were employed to assess a generated text based on surface-level similarities, measured through the n-gram overlap between the generated and reference texts.
Recent trends in summarization evaluation lean towards semantic-level assessments, moving beyond direct word overlap comparisons. Notable metrics embracing this approach include BERTScore Zhang et al. (2020), MoverScore Zhao et al. (2019), BARTScore Yuan et al. (2021), BLEURT Sellam et al. (2020), and variations thereof. By leveraging pretrained language models, these metrics focus on capturing semantic content, providing a more nuanced and accurate evaluation of summarization system outputs.
@@
With the widespread use of generation models across diverse domains, the need for reference-free evaluation metrics has surged. In response to this challenge, recent attention has been directed towards metrics that enable the evaluation of generated texts solely based on source documents, especially when annotated reference texts may not be available for new domains Böhm et al. (2019); Gao et al. (2020); Wu et al. (2020); Chen et al. (2021); Scialom et al. (2021); Honovich et al. (2021); Zhong et al. (2022); Liu et al. (2023).
@@
Despite their widespread use, automatic evaluation metrics have notable shortcomings. These metrics may not be robust when faced with challenges such as spurious correlations, noise, or out-of-domain texts (Sai et al., 2021; Vu et al., 2022; Durmus et al., 2022; Zhao et al., 2023; He et al., 2023). Furthermore, their effectiveness diminishes when evaluating very long documents (Amplayo et al., 2022). There is also evidence suggesting a potential bias towards ranking extractive summaries higher than abstractive ones Amplayo et al. (2022).
Traditional reference-based evaluation metrics such as ROUGE or BLEU have been criticized for their inability to measure content quality or capture syntactic errors Reiter and Belz (2009). Consequently, these traditional metrics often exhibit weak correlations with human judgements, demonstrating that they cannot accurately reflect the real-world performance of generation systems Peyrard (2019); Mathur et al. (2020). For example, they might assign high scores to outputs that are fluent but meaningless and unfaithful, as long as many of the same words are used Gehrmann et al. (2021). Although embedding-based metrics (e.g., BERTScore) show improved performance in similarity measurement, they are still inadequate for assessing the extent of shared information between two summaries, a crucial indicator of summary information quality Deutsch and Roth (2021).
Reference-free metrics, on the other hand, exhibit a bias towards outputs generated by models that are more similar to their own Deutsch et al. (2022). To the best of our knowledge, this study represents the initial attempt to perform an exploration, which has not yet been undertaken systematically. Additionally, question-answering-based reference-free metrics for summarization evaluation are prone to inheriting errors within summaries Kamoi et al. (2023).
Metrics based on Large Language Models, which are capable of conducting both reference-based and reference-free evaluations, typically demonstrate superior correlations with human quality judgements across diverse NLG tasks and evaluation dimensions Deutsch et al. (2022). While prior work has reported that LLM-based metrics prefer LLM-generated text, raising a concern about the shortcomings of LLMs as evaluators Liu et al. (2023), our work conducts a systematic evaluation to address a fundamental question: Do language model-driven evaluation metrics inherently display bias favouring texts generated by the same underlying language model? We explore this question across both reference-based and reference-free evaluations and for a range of different large language models.
@@ 3 Methodology
To investigate the impact of the model’s self-bias—determining whether a language model-based evaluator favours outputs generated by a similar language model—we conduct a comprehensive series of experiments involving both quantitative comparisons and qualitative analysis. Our quantitative comparisons involve using language models of varying sizes and finetuning configurations as both the evaluator and generator models. This structured approach enables us to systematically examine the potential bias across different LM configurations. Subsequently, we verify the results through qualitative analysis using a subset of models’ summaries that are accompanied by human evaluation to further demonstrate that higher scores produced by evaluators as a result of self-bias do not necessarily correlate with higher quality generated outputs.
@@ 3.1 Evaluators
We describe the evaluation process as follows: given a source text s, a human written reference r, generate a hypothesis h, which can be represented as:
@@
where h denotes hypothesis, a refers to the aspect to evaluate, and $\mathcal{S}$ denotes supplementary text (i.e., s or r) that is employed alongside evaluations in various settings Fu et al. (2023). For instance, it could be the source text s in a reference-free scenario which assesses the summary based on the source article directly Fabbri et al. (2021). Whereas in the reference-based paradigm, the evaluation considers semantic overlap between the generated hypothesis h (e.g. model generated summaries) and reference summaries r Bhandari et al. (2020).
The evaluators (i.e. based on BART, T5, GPT model variants as well as Cohere) utilised in our study all share a conditional probability paradigm, which can generally be formulated as
$\!\!\!\!\!\!\!\!\mathrm{Score}(\textbf{{h}}|d,a,\mathcal{S})=\sum_{t=1}^{m}w_{t}\log{p}(h_{t}|\textbf{{h}}_{<t},\mathcal{S},\theta).$ (2)
Here $\theta$ is the model parameter, $d$ refers to the task description and $w_{t}$ denotes the weight of the token at position $t$, where previous works normally treat each token equally Yuan et al. (2021); Fu et al. (2023). We provide further descriptions of each type of evaluator below.
@@
Leveraging generative models to conduct evaluation has been further advanced with various of more recent large language models Fu et al. (2023), showing a great performance and covering a rich variety of aspects for comprehensive evaluations. With an understanding of natural language instructions, GPTScore (including GPT-X and FLAN-T5 models) can perform intricate and personalized assessments without additional training.
@@
We additionally include Cohere, the more recent language model to enrich our assessments. The evaluation scores assigned by the model is calculated according to Eq. 2, aligned with BARTScore, T5Score, and GPTScore.
@@ 3.2 Generation Models
To ensure the reproducibility of our analysis, we exclusively employ publicly available checkpoints for the utilized models. Apart from the GPT3-Curie model that is taken from the OpenAI API and generation model obtained from Cohere, the rest of the models are taken from the Hugging Face model hub [^3] https://huggingface.co/models..
@@ 3.3 Datasets
The SummEval benchmark contains summaries generated by various summarization models (i.e. BART, T5 and GPT2) for 100 articles from the CNN/DM test set, with each summary supplemented by human annotations. More specifically, SummEval incorporates human annotations by both expert and crowd-sourced human annotators, targeting dimensions of coherence, consistency, fluency, and relevance. Ratings are on a scale of 0 to 5, with higher values indicating better performance.
Similarly, RoSE contains summaries generated by recent generative models based on CNN/DM documents, accompanied by their corresponding human evaluations. We use 100 summaries from each of the BART and GPT-3 models from the ROSE benchmark. The RoSE benchmark proposed an assessment protocol termed “Atomic Content Units” (ACUs) Liu et al. (2022). ACU score gauges quality of evaluated summaries based on whether the presence of single facts (i.e., atomic facts) from reference are included in the evaluated summaries. ACU score is calculated by ACU matching:
@@
where $\mathcal{A}$ is a set of ACUs from gold summaries and $\mathcal{A}_{s}$ denotes the ACUs of candidate summary s.
@@
SummEval-BART 5.00 2.67 4.57 4.67
SummEval-T5 5.00 2.33 4.52 4.67
SummEval-GPT2 5.00 1.33 3.57 3.58
Table 1: Distribution of human annotation scores on the RoSE and SummEval datasets, where in RoSE we consider the ‘ACU’ score, and in SummEval we focus on four aspects—‘Coherence’, ‘Consistency’, ‘Fluency’, and ‘Relevance’—as evaluated by expert annotators. The scores for SummEval are obtained by averaging the scores across all aspects and evaluations from all annotators.
@@
Figure 2: Assessing Bias in the CNN/DM Dataset using heatmaps in the reference-free setting. Observing the darkest cells along the diagonal line, from the top left to the bottom right, indicates a distinct bias among evaluators towards their respective models. All evaluator scores are normalized to a range between 0 and 1. Additionally, the number in the bracket represents the average length of summaries (measured in words) produced by the respective model.
@@ 3.4 Quantitative Comparisons
Due to the nature of log probabilities, original scores from each evaluator is be negative, and a higher score indicates better quality according to the evaluator. When weights $w_{t}$ in Eq. 2 are treated equally, the evaluation protocols of BARTScore, T5Score, and GPTScore are all conditional probability paradigms. To ensure comparability among the scores provided by 20 distinguished evaluators, a uniform normalization process is applied to the scores generated by each evaluator. The normalization procedure standardizes the scores across a scale ranging from 0 to $\alpha$  [^6] In our work, we set parameter $\alpha$ to 1. as formulated in Eq. 4, where $X_{i,j}$ indicates scores evaluated by the $j$-th evaluator on summaries generated by the $i$-th generator.
$X_{i,j}^{norm}=\frac{\alpha(X_{i,j}-\min_{i}X_{i,j})}{\max_{i}X_{i,j}-\min_{i}X_{i,j}}$ (4)
In this context, a normalized score of $\alpha$ signifies the highest quality attributed by the evaluator, while a score of 0 indicates the lowest quality.
As the length of the generated summary is a key factor influencing the evaluation results, we further analyse the impact of lengths for the content generated by the models along with the experiments. In this regard, we also compute the correlations between the length of the text and the scores assigned by evaluators to identify trends in evaluators’ preferences.
@@ 3.5 Qualitative Analysis
For qualitative analysis, we employ Spearman Correlations Zar (2014) and Kendall Correlations Freedman et al. (2007), which respectively assess monotonic relationships and order associations between human evaluations and LM evaluator scores. They are common metrics for assessing correlations with human judgements.
For the SummEval dataset, we calculate the correlations for four aspects (i.e. Coherence, Consistency, Fluency and Relevance ), aligned with the reference-free input setting in the evaluation protocol as specified by Yuan et al. (2021). For the evaluations based on the RoSE benchmark, we use ACU annotations that are suited for reference-based summary salience evaluation. Therefore, we employ the correlation values obtained from the SummEval dataset for the reference-free setting and those from the RoSE benchmark for the reference-based setting.
@@ 4 Experimental Results
4.1 Quantitative Comparisons: Assessing Self-Bias in LM-Evaluators Towards Their Own Output
Figure 3: Assessing Bias on CNN/DM Dataset using heatmaps in the reference-based setting. Observing darker cells along the diagonal line indicates potential self-bias. All evaluator scores are normalized to a range between 0 and 1. Additionally, the number in the bracket represents the average length of summaries (measured in words) produced by the respective model.
Figures 2 and 3 display heatmaps presenting evaluator scores for various summaries generated by different generators from CNN/DM documents in reference-free and reference-base settings, respectively. These scores are computed by averaging the individual scores of the selected 500 documents. In both heatmaps, we observe darker cells along the diagonal line, running from the top left to the bottom right. This indicates the potential evaluator bias towards their corresponding generator models i.e., self-bias. However, this bias is notably more pronounced in the reference-free setting, commonly used for factuality evaluation Koh et al. (2022).
Furthermore, as shown in Figure 2, we note a distinct trend: T5-based generators, whether fine-tuned or not, tend to receive higher scores when assessed using different T5Score variations compared to evaluations using BARTScore, GPTScore, or Cohere. This results in a concentrated dark rectangle at the heatmap’s centre. Similarly, we observe a parallel trend for BART-based generators, whether fine-tuned or not.
Meanwhile, evaluators tend to assign higher ranks to generators trained on the same dataset as themselves, rather than to those fine-tuned on different datasets (see Figure 6). For example, when using T5 models fine-tuned on the XSUM dataset as evaluators, there is a noticeable preference for BART-XSUM generators over T5-vanilla models, even though the evaluations are performed for the CNN Daily dataset. We observe the same pattern on summaries generated based XSUM documents.
Figure 4: Heatmaps of evaluation scores on the SummEval & RoSE benchmarks for the reference-free and reference-based setting. We use the reference-free setting for SummEval and the reference-based setting for RoSE, aligning with the specific aspects each benchmark emphasizes.